# Alchemist: Towards the Design of Efficient Online Continual Learning System

### v1

@@ 4.1.1. Recording Activations
A naïve design for reusing the activations of the model generated responses is to record and save all the activations until the end of generation during serving (i.e., torch.enable_grad for both prefill and decode phases). However, this could introduce prohibitive overhead. As aforementioned, enabling activation recording could incur up to 35% latency overhead in the decoding phase. If we enable it for both prefill and decode phases, this overhead would apply to every forward pass, eventually leading to 35% latency overhead end-to-end. On the other hand, completely disabling activations recording for both prefill and decode phase to maintain serving latency eliminates the possibility of reusing activations calculated from serving. The loss calculation then would require full prefill on the prompt with chosen and rejected responses twice, leading to redundant computation.
@@
Nevertheless, a balanced approach is possible: we can enable activation recording for only the prefill phase, and save the KV cache and the activations associated with it. This design allows loss calculation to only attend to the response texts, skipping the prompts. The overhead from recording activations then only applies to the prefill phase and will be amortized as generation proceeds.
@@
Since both the chosen response and the rejected chosen share the same prompt, the two forward passes can share the KV cache as well, further reducing redundant computation. Besides the computation saving, sharing KV cache also means sharing the activations associated with it among the two forward passes, hence also reducing unnecessary duplication in memory.
@@ 4.2. Schedule activation reuse
To address this issue, Alchemist employs a straightforward yet effective strategy: it preempts the training job and immediately switches to the serving job as soon as a serving query arrives. Leveraging the hook functionality provided by frameworks such as PyTorch, Alchemist injects preemption hooks at the start of each layer’s backward function call. These hooks check for any serving forward pass that are currently running or queued. If so, the backward pass is paused until the serving forward operations have completed and the queue is empty.
In the case like DPO loss calculation, where both backward pass and forward pass are required during training, similarly preemption hooks will be injected into each layer’s forwards. Through this approach, Alchemist minimizes the overlap between training and serving processes, effectively minimizing the impact on serving latency from the training side. Together with the design in §4.1.1, Alchemist could fulfill the latency and activation reuse requirement from both R1 and R2.
@@ 4.3. Alchemist activation offloader
However, saving all the activations can risk out-of-memory (OOM). As the context length grows, even if only saving LoRA adapter’s activations and de-duplicate activations with input prompt KV cache reuse as described in §4.1.1, the activations from both the prefill phase and loss calculation can easily reach 40 GB as shown in Figure 6 when total trained token (i.e., both prompt and response) is 3000 tokens. If we save all these activations while serving at the same time, such large memory overhead can severely limit the context length or the batch size for the serving side.
Figure 6. Peak memory usage with DPO. Even if leveraging the input prompt’s KV cache as described in §4.1.1, naively saving all required activations in GPU memory on the train side can still introduce significant memory overhead. This could severely limit the supported context length or batch size on the serving side before OOM. §4.3.
Naively offloading every activation to CPU memory, on the other hand, will leave GPU memory underutilized since serving queries does not necessarily always require the full memory space. When offloading more activation as we need, we are actually introducing unnecessary loading overhead when we load activations that could have been kept on GPU memory back to the GPU.
Hence, the goal of Alchemist ’s offloader is to offload activations as needed rather than offloading every single activation to CPU memory. To achieve this, Alchemist asynchronously copy the activations during the forward pass to a pre-allocated and pinned CPU memory. When we determine a serving query requires more GPU memory, we dynamically free activations until there is enough memory to serve the incoming queries.
@@ 4.3.1. Offloading map
Hence, before runtime, with a given model and GPU, we can profile a mapping, namely offloading map, with these three parameters as input and the number of bytes to free as the output. In the real system, for the system’s simplicity, we restricted the minimum unit for freeing to one layer’s activations, instead of bytes or activations. By referring to this profile mapping, when we start each serving forward at runtime, we may know how many layers’ activations we must free to avoid the risk of OOM while serving.
@@ 4.3.2. Pipelining
Since backpropagation proceeds from the output layer to the input layer, the optimal strategy to minimize the wait time is to prioritize retaining the activations of the higher layers (i.e., layers close to the output layer). In another word, Alchemist should free the activations in the order of forward pass (i.e., from input layer to the output layer). This ensures that, when backpropagation begins, the gradients for the higher layers can be computed immediately, as their activations remain in GPU memory. Meanwhile, as backpropagation operates on the layers with activations ready in GPU memory, we can pipeline and prefetch lower layers’ (i.e., layers closest to the input layer) activations.
@@ 5. Implementation
Following the design depicted in §4, we implemented a prototype of Alchemist with 1.5K LOC in Python. We rely on commonly used third-party libraries, like PyTorch and Transformers in Alchemist’s implementation. To ensure Alchemist’s accessibility and usability, instead of providing as a standalone serving engine, we implement Alchemist as a plugin that can be integrated into existing state-of-the-art serving engines like vLLM (Kwon et al., 2023), SGLang (Zheng et al., 2024), et al..
Alchemist cache class:  Different training methods as well as different loss functions may require different inputs that we can cache from serving. For example, with continual pretrain with cross entropy loss, the loss function only requires the prefill output and prompt text, but for continual preference alignment with DPO loss, it would require prefill phase KV cache, the accepted response and rejected response text. To make Alchemist extensible to all possible training methods and loss functions, we provide AlchemistCacheBase abstract class, allowing users to wrap different kinds of cache they would like to store from serving. In the class, we also implement synchronization functions indicating the label, if needed, for the corresponding cache is ready, for instance, is_ready() or wait_ready().
@@
    push(cache: AlchemistCacheBase): Within the serving loop, users call this function to push the calculated activations and other values wrapped in AlchemistCacheBase.
@@
    push_label(label: str): If label is required, users call this function to push the label to Alchemist
@@
    pull() -> AlchemistCacheBase: When implementing customized training methods and loss functions, users use this function to pull the cached activations and other values that required by the training method and loss calculation (e.g., prompt text, label, et al.). This function will be blocking as it calls AlchemistCacheBase’s wait_ready() to wait until label is ready.
@@
With this implementation, besides users’ customized training implementation, the users of Alchemist are only expected to make minimal changes to their existing serving engines and model implementations..
Alchemist offloader:  As a model inheriting from Alchemist, we also register PyTorch’s saved_tensor_hook (PyTorch Tutorials, 2025) to each layer of the models. This hook function is called each time an activation is recorded by PyTorch’s autograd framework. We use this hook to (1) asynchronously copy the activations from GPU memory to CPU memory as described in §4.3 and (2) record the ownership of the activations (i.e., which layer’s backward requires this activation). This ensures we free the correct activations when certain layers’ activations are required to be freed. We use a CUDA Stream differed from PyTorch’s default stream to overlap the computation and data movement between GPU and CPU (PyTorch Contributors, 2025b). We primarily rely on PyTorch’s cache allocator (PyTorch Contributors, 2025a) to virtually free the memory instead of physically de-allocate activations’ underlying buffer from GPU memory. This allows serving job to immediately reuse the underlying buffer by overwriting the value in the buffer as needed.
@@ 6. Evaluation
We evaluate Alchemist’s performance by comparing it with a separate training cluster that does not reuse serving activations, which serves as our baseline. We use Llama-3.1-8B (Touvron et al., 2023) with a LoRA adapter of rank 8 as the example training and serving model throughout the evaluation process. The evaluation testbed is equipped with a single Nvidia A100 80GB SXM GPU (NVIDIA, 2025) and two AMD EPYC 7V12 64-Core CPUs (AMD, 2025), along with a total of 1.73 TB of CPU memory. As offloading may require a relatively large amount of memory, we use taskset (Love, 2025) to set the evaluation process affinity to one NUMA node at all times to avoid possible NUMA effects.
We implement a naïve serving engine for fast prototyping and easy integration with Alchemist. We use ShareGPT (anon8231489123, 2025) as our evaluation dataset. Since ShareGPT does not provide timestamps, and following the approach used in many prior works (Kwon et al., 2023; Zheng et al., 2024; Miao et al., 2024a), we generate serving traces by sampling from a Poisson distribution when evaluating at varying request rates.
@@ 6.1. Training throughput improvement
We first evaluate Alchemist ’s training throughput across different token lengths (see Fig. 8). When performing continual pretraining (CPT) with cross-entropy loss (upper figure), Alchemist (in blue) consistently outperforms the baseline (in red), delivering over a 1.7x improvement in throughput regardless of token length. For continual preference alignment (CPT) with DPO loss, Alchemist achieves between 1.26x and 1.68x improvement, depending on the token length.
@@
Moreover, while the baseline runs out of memory (OOM) for token lengths exceeding 3000 tokens (marked in red “x”), Alchemist can support up to 7000 tokens, thanks to the memory savings from reusing the KV cache.
@@ 6.2. Alchemist’s impact on serving latency
We evaluate Alchemist’s serving latency overhead by measuring the Time-per-Token (TPT) during serving as the metric. We start by measuring the TPT distribution with under the average request rate of 1.7 query per second (QPS). As the left figure in Fig. 9, Alchemist poses only minimal impact on the TPT distribution and does not introduce large tail latency during serving.
We further evaluate Alchemist’s impact on the average TPT at various request rates, the result, shown in the right figure in Fig 9, demonstrate that Alchemist increases the average TPT by at most 3%.
During the attempt to further reducing such overhead, we observe the bulk of the overhead actually come from the PyTorch’s invocation on the registered hook functions. In another word, even if we register a empty hook function, such overhead would still incur. If we are willing to sacrifice the usability of Alchemist and add the currently used hook functions as part of models’ implementation directly, we can further reduce the latency overhead.
@@ 6.3. Memory saving with KV reuse
In the case of continual preference alignment, Alchemist leverages serving prompt’s KV cache reuse reduce the memory duplication of the activations. To evaluate the memory saving, we start by compare the max token length before OOM. As illustrated in the left figure of Fig. 10, without reusing input prompt’s KV cache, baseline can only support up to 3000 tokens with our hardware setup. Meanwhile, with the same setup, by remove the duplication in memory with KV cache reuse, Alchemist can support up to 7000 tokens.
@@
6.4. Offloading’s impact
Figure 11. Impact of offloading activations. Alchemist’s improvement drops as we offloading more activations to CPU memory. This is expected since more activations offloaded means Alchemist must load more activations back from CPU memory to GPU memory which can be expensive. The normalized improvement eventually reaches 1x meaning that, based on the hedging map described in §4.3.3, Alchemist decides it is more efficient to recompute the forward passes instead of loading the activations back from CPU memory. §6.4.
When offloading is required to accommodate incoming queries’ memory requirement, it unavoidably will slow down the training procedure because of the delay when loading the activation back from CPU memory to GPU memory. To understand the impact of offloading activations, we vary ratio of offloaded activations and measure the training throughput.
As Fig. 11 shown, the normalized improvement decreases as we offloading more activations but eventually bottomed at 1x when offloading ratio is near 30%. This is because after querying the hedge map described in §4.3.3, Alchemist decides it would be more efficient to recompute the forward pass instead of load the activations back from the CPU memory, whose computation is the same as the baseline’s.
@@ 7. Discussion
Parallelism Strategy Across Training and Serving Processes:  A possible concern is that different parallelism strategies between serving and training may prevent the reuse of serving activations in the training process. We leave this as future work.

### v2

@@ 4.1. Record and cache activation
As aforementioned, enabling activation recording could incur up to 35% latency overhead in the decoding phase. If we enable it for both prefill and decode phases, this overhead would apply to every forward pass, eventually leading to 35% latency overhead end-to-end. On the other hand, completely disabling activations recording for both prefill and decode phase to maintain serving latency eliminates the possibility of reusing activations calculated from serving. The loss calculation then would require full prefill on the prompt concatenated with chosen and rejected responses twice, leading to redundant computation.
@@
Nevertheless, a balanced approach is possible: we can enable activation recording for only the prefill phase, and save the KV cache and the activations associated with it. This design allows loss calculation to only attend to the response texts, skipping the prompts and reusing part of the activations calculated during serving. Similar to the case of continual pretrain, such overhead then only applies to the prefill phase and will be amortized as generation proceeds.
@@
Since both the chosen response and the rejected response share the same prompt, the two forward passes can share the KV cache as well, further reducing redundant computation. Besides the computation saving, sharing KV cache also means sharing the activations associated with it among the two forward passes, hence also reducing unnecessary duplication in memory.
@@ 4.2. Schedule activation reuse
To address this issue, Alchemist employs a straightforward yet effective strategy: it preempts the training job and immediately switches to the serving job as soon as a serving query arrives, temporally sharing GPU resources among serving and training jobs. Leveraging the hook functionality provided by frameworks such as PyTorch, Alchemist injects preemption hooks at the start of each layer’s backward function call. These hooks check for any serving forward pass that is currently running or queued. If so, the backward pass is paused until the serving forward operations have completed and the queue is empty.
In the case like DPO loss calculation, where both backward passes and forward passes are required during training, similarly, preemption hooks will be injected into each layer’s forward. Through this approach, Alchemist minimizes the overlap between training and serving processes, effectively minimizing the impact on serving latency from the training side. Together with the design in §4.1, Alchemist could fulfill the latency and activation reuse requirement from both R1 and R2.
@@ 4.3. Alchemist activation offloader
However, saving all the activations can risk out-of-memory (OOM). As the context length grows, even if only saving LoRA adapter’s activations and de-duplicate activations by sharing input prompt KV cache as described in §4.1, the activations from both the prefill phase and DPO loss calculation can easily reach 40 GB as shown in Figure 6 when total trained token (i.e., both prompt and responses) is 3000 tokens. If we save all these activations while serving at the same time, such large memory overhead can severely limit the context length or the batch size for the serving side.
Figure 6. Peak memory usage with DPO loss. Even if leveraging the input prompt’s KV cache as described in §4.1, naively saving all required activations in GPU memory on the train side can still introduce significant memory overhead. This could severely limit the supported context length or batch size on the serving side before OOM. §4.3.
Naively offloading every activation to CPU memory, on the other hand, will leave GPU memory underutilized since serving queries does not necessarily always require the full memory space. When offloading more activations than we actually need, we introduce unnecessary loading overhead when we load the activations that could have been kept on GPU memory back to the GPU.
Hence, the goal of Alchemist ’s offloader is to offload activations as needed rather than offloading every single activation to CPU memory. To achieve this, Alchemist asynchronously copy the activations during the forward pass to a pre-allocated and pinned CPU memory. Unless, even if without overlapping with serving jobs, the activations themselves are too large to be fitted, we will not immediately free the copied activations from GPU memory. Only when we determine a serving query requires more GPU memory, we then dynamically free activations in GPU memory until there is enough memory to serve the incoming queries.
@@ 4.3.1. Offloading map
Hence, before runtime, with a given model and GPU, we can profile a mapping, namely offloading map, with these three parameters as input and the number of bytes to free as the output. In the real system, for the system’s simplicity, we restricted the minimum unit for freeing to one layer’s activations, instead of bytes or activations. By referring to this profile mapping, when we start each serving forward at runtime, we may know how many layers’ activations we must free to avoid the risk of OOM while serving. In addition, we may also know if there will be OOM during activations recording due to the size of activations itself. In this case, we directly offload the activations to CPU memory without retaining them in GPU memory at all.
@@ 4.3.2. Pipelining
Since backpropagation proceeds from the output layer to the input layer, the optimal strategy to minimize the wait time is to prioritize retaining the activations of the higher layers (i.e., layers close to the output layer). In other words, Alchemist should free the activations in the order of forward pass (i.e., from input layer to the output layer). This ensures that, when backpropagation begins, the gradients for the higher layers can be computed immediately, as their activations remain in GPU memory. Meanwhile, as backpropagation operates on the layers with activations ready in GPU memory, we can pipeline and prefetch lower layers’ (i.e., layers closest to the input layer) activations.
@@ 4.3.3. Hedging map
Note here, for loss functions that require multiple forward passes with the same prompt, even if Alchemist needs to recompute the forward passes, it still leverages the KV cache sharing technique introduced in §4.1. This means that it will prefill the user prompts only once and share the corresponding KV cache across multiple forward passes to accelerate the recomputation process and reduce the memory footprint.
@@ 5. Implementation
Following the design depicted in §4, we implemented a prototype of Alchemist with 1.5K LOC in Python. We rely on commonly used third-party libraries, like PyTorch and Hugging Face’s transformers (Wolf et al., 2020) in Alchemist’s implementation. To ensure Alchemist’s accessibility and usability, instead of providing it as a standalone serving engine, we implement Alchemist as a plugin that can be integrated into existing state-of-the-art serving engines like vLLM (Kwon et al., 2023), SGLang (Zheng et al., 2024), et al..
Alchemist cache class:  Different training methods as well as different loss functions may require different inputs that we can cache from serving. For example, with continual pretrain using cross-entropy loss, the loss function only requires the prefill output and prompt text, but for continual preference alignment with DPO loss, it would require prefill phase KV cache, the accepted response, and rejected response text. To make Alchemist extensible to all possible training methods and loss functions, we provide AlchemistCacheBase abstract class, allowing users to wrap different kinds of cache they would like to store from serving. In the class, we also implement synchronization functions indicating the label, if needed, for the corresponding cache is ready, for instance, is_ready() or wait_ready().
@@
    push(cache: AlchemistCacheBase): Within the serving loop, users call this function to push the serving activations and other values wrapped in AlchemistCacheBase.
@@
    push_label(label: str): If labels are required, users call this function to push the label to Alchemist.
@@
    pull() -> AlchemistCacheBase: When implementing customized training methods and loss functions, users use this function to pull the cached activations and other values that are required by the training method and loss calculation (e.g., prompt text, label, et al.). This function is blocking as it calls AlchemistCacheBase’s wait_ready() to wait until the label is ready when it is required.
@@
With this implementation, besides users’ customized training implementation, the users of Alchemist are only expected to make minimal changes to their existing serving engines and model implementations.
Alchemist offloader:  As a model inheriting from Alchemist, we also register PyTorch’s saved_tensor_hook (PyTorch Tutorials, 2025) to each layer of the models. This hook function is called each time an activation is recorded by PyTorch’s autograd framework. We use this hook to (1) asynchronously copy the activations from GPU memory to CPU memory as described in §4.3 and (2) record the ownership of the activations (i.e., which layer’s backward requires this activation). This ensures we free the correct activations when certain layers’ activations are required to be freed. We use a CUDA Stream different from PyTorch’s default stream to overlap the computation and data movement between GPU and CPU (PyTorch Contributors, 2025b). We primarily rely on PyTorch’s cache allocator (PyTorch Contributors, 2025a) to virtually free the memory instead of physically de-allocating activations’ underlying buffer from GPU memory. This allows the serving job to immediately reuse the underlying buffer by overwriting the value in the buffer as needed.
@@ 6. Evaluation
Baseline:  We evaluate Alchemist’s performance by comparing it with a separate training cluster that does not reuse serving activations nor share the prompt’s KV cache, which serves as our baseline.
Hardware:  The evaluation testbed is equipped with an Nvidia A100 80GB SXM GPU (NVIDIA, 2025) and two AMD EPYC 7V12 64-Core CPUs (AMD, 2025), along with a total of 1.73 TB of CPU memory. As offloading may require a relatively large amount of memory, we use taskset (Love, 2025) to set the evaluation process affinity to one NUMA node to avoid possible NUMA effects.
Models:  We use Llama-3.1-8B (Touvron et al., 2023), Mistral-v0.2-Instruct (Jiang et al., 2023), and Phi-4 (Abdin et al., 2024) as our evaluation models.
Dataset:  We randomly sampled the prompts from ShareGPT (anon8231489123, 2025), which were collected from real users’ conversations with ChatGPT, as our input dataset. We set the output length to be 128 tokens for all queries. In the case where we evaluate Alchemist under various serving loads (i.e., varied queries-per-second, QPS), since ShareGPT does not provide timestamps, following the approach used in many prior works (Kwon et al., 2023; Zheng et al., 2024; Miao et al., 2024a), we generate traces by sampling from a Poisson distribution at varying request rates.
Serving:  We implement a naïve serving engine for fast prototyping and integration with Alchemist using Hugging Face.
Training:  For continual pretrain, we use CrossEntropy loss as an example. For continual preference alignment, we use DPO loss as an example training method. For DPO loss, we assume model-generated responses are the rejected responses and generate a random tensor with the same length as the chosen response. We generate the chosen one with minimal delay after the model finishes generating to mimic users providing feedback immediately once they see the model output or auto-labeling LLM. Models are attached with a LoRA adapter of rank 8 as the example training settings for the evaluation process.
Memory profiler settings:  For Llama-3.1-8B and Mistral-v0.2-Instruct, we use a step size of 500 tokens for cached token length and incoming query token length and a batch size of 5 when profiling memory consumption. For Phi-4, due to its larger model size, to reduce the profiling error, we decrement the step size to 250 tokens for cached and incoming query.
Metrics:  We use token-per-second trained to evaluate the training throughput. Since the time Alchemist can start its training jobs depending on the actual serving workload, to isolate this factor and focus on Alchemist’s sole efficiency, we only count the serving idle time when calculating training throughput. When evaluating the serving side, we use time-per-token (TPT) output as the metric.
@@ 6.1. Training throughput improvement
We first evaluate Alchemist ’s training throughput when no offload is needed (i.e., serving workload is light). We sample various lengths of input tokens from the ShareGPT dataset to understand Alchemist’s performance under various token lengths. As shown in Figure 8), when performing continual pretraining (CPT) with cross-entropy loss (upper figure), Alchemist (in blue) consistently outperforms the baseline (in red), delivering over a 1.72x improvement in throughput regardless of token length. For continual preference alignment (CPT) with DPO loss, Alchemist achieves between 1.26x and 1.68x improvement, depending on the token length.
@@
Moreover, while the baseline runs out of memory (OOM) for token lengths exceeding 3000 tokens for Llama-3.1-8B and Mistral and 1500 tokens for Phi-4(marked in red “x”), Alchemist can support up to 7000 tokens and 2500 tokens respectively, thanks to the memory savings from reusing the KV cache.
6.2. Offloading’s impact
Figure 9. Impact of offloading activations with DPO loss. Alchemist’s improvement drops as we increase QPS requiring Alchemist to offload more activations to CPU memory due to the larger serving batch size and corresponded higher memory requirements. This is expected since more activations offloaded means Alchemist must load more activations back from CPU memory to GPU memory which can be expensive.
Next, we evaluate Alchemist’s performance when offloading is required to accommodate incoming queries’ memory requirement due to either larger QPS or longer input token length. We use Alchemist with DPO loss as the evaluation example since DPO loss requires two forward passes hence having a larger memory footprint. We increase the QPS until there is no idle time for Alchemist to run training jobs. To better illustrate the impact of offloading, when sampling prompts from ShareGPT dataset, we set a minimal token length of 4000 for Llama-3.1-8B and Mistral-v0.2-Instruct and 2500 for Phi4. These numbers were chosen as they are the minimum token length we started to observe the impact of offloading under aforementioned QPS and training settings.
As Figure 9 shows, as we increase the QPS, due to the larger batch size during serving, Alchemist is required to offload more layers to CPU memory. This, as expected, will reduce the training throughput because of the delay when loading the activation back from CPU memory to GPU memory. Especially, in the case of Phi-4, due to its larger model size with higher memory requirements, when at 1.7 QPS, Alchemist decides to recompute the majority of the forward passes after querying its hedge mapping as described in §4.3.3.
Since, during recomputation, Alchemist still only prefills the input prompt once and shares the KV cache across two responses, it can significantly reduce the memory footprint and avoid OOM, unlike the baseline. Hence, even with offloading, Alchemist is still better than the baseline.
@@ 6.3. Alchemist’s impact on serving latency
To understand Alchemist’s impact on the serving side, we start by measuring the TPT distribution under the average request rate of 1.7 queries per second (QPS) with Llama-3.1-8B and DPO loss. As the left figure in Fig. 10, Alchemist poses only minimal impact on the TPT distribution and does not introduce large tail latency during serving.
We further evaluate Alchemist’s impact on the average TPT at various request rates. The result, shown in the right figure in Fig 10, demonstrates that Alchemist increases the average TPT by at most 3%.
During the attempt to further reduce such overhead, we observe the bulk of the overhead actually comes from the PyTorch’s invocation on the registered hook functions. In other words, even if we register an empty hook function, such overhead would still incur. If we are willing to sacrifice the usability of Alchemist and add the currently used hook functions as part of models’ implementation directly, we can further reduce the latency overhead.
@@ 6.4. Memory saving with KV reuse
In the case of continual preference alignment with DPO loss and Llama-3.1-8B model, Alchemist leverages the serving prompt’s KV cache reuse to reduce the memory duplication of the activations. To evaluate the memory saving, we start by comparing the max token length before OOM. As illustrated in the left figure of Fig. 11, without reusing the input prompt’s KV cache, baseline can only support up to 3000 tokens with our hardware setup. Meanwhile, with the same setup, by removing the duplication in memory with KV cache reuse, Alchemist can support up to 7000 tokens.
@@ 7. Discussion
Parallelism Strategy Across Training and Serving Processes:  A possible concern is that different optimal parallelism strategies between serving and training may impact the training throughput when co-locating serving and training. We leave this as future work.